OpenStudy (anonymous):

Tyler is twice as old as his son. Ten years ago, Tyler was three times as old as his son. How old are they now?

OpenStudy (anonymous):

I understand the equation is T=2s and t-10=3(5-10) but do I just plug in my own numbers from here?

OpenStudy (anonymous):

please explain how you went from t-10= 3s-30 to the next step

OpenStudy (anonymous):

got it, so then you wind up with 2s=3s-20 what did you do to get s

OpenStudy (anonymous):

i wish

OpenStudy (anonymous):

ok i got it thanks